# Symbiosis: Multi-Adapter Inference and Fine-Tuning

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) allows model builders to capture the task-specific parameters into adapters, which are a fraction of the size of the original base model. Popularity of PEFT technique for fine-tuning has led to the creation of a large number of adapters for popular Large Language Models (LLMs). However, existing frameworks fall short in supporting inference or fine-tuning with multiple adapters in the following ways. 1) For fine-tuning, each job needs to deploy its dedicated base model instance, which results in excessive GPU memory consumption and poor GPU utilization. 2) While popular inference platforms can serve multiple PEFT adapters, they do not allow independent resource management or mixing of different PEFT methods. 3) They cannot share resources (such as base model instance) between inference and fine-tuning jobs. 4) They do not provide privacy to users who may not wish to expose their fine-tuned parameters to service providers. In Symbiosis, we address the above problems by enabling the as-a-service deployment of the base model. The base model layers can be shared across multiple inference or fine-tuning processes. Our split-execution technique decouples the execution of client-specific adapters and layers from the frozen base model layers offering them flexibility to manage their resources, to select their fine-tuning method, to achieve their performance goals. Our approach is transparent to models and works out-of-the-box for most models in the transformers library. Our evaluation on Llama2-13B shows that compared to baseline, Symbiosis can fine-tune 4X more adapters on the same set of GPUs in the same amount of time.
1 Introduction
@@
Parameter Efficient Fine Tuning (PEFT) [23] is a popular method to fine-tune the existing foundation models with various fine-tuning methods, e.g., LoRA, IA3, and AdaLoRA. The adapters created using PEFT are a fraction of the size of the base model used, often consuming only 10s of MBs of accelerator memory, as opposed to the base model that consumes several Gigabytes. The resulting efficiency and low cost in fine-tuning adapters have led to the creation of hundreds of adapters for many popular models, such as Llama, Gemma, StarCoder, etc. For instance, an LLM agentic system may fine-tune and serve several adapters that address different use cases. Or, they may want to evaluate a range of adapter hyperparameters, e.g., LoRA rank, alpha, etc. to reach the best possible model accuracy.
The problem with the existing platform that allows users to serve or fine-tune multiple PEFT adapters is that they do not fully utilize GPU’s computational capability [27, 31]. For instance, to fine-tune multiple PEFT adapters of a model at once, separate jobs must be launched in parallel, each needing a dedicated model instance. The model instances often occupy the GPU’s memory well before fully exploiting its computational capability. Moreover, even when the model instances can be accommodated in GPU memory, they may not leave enough free memory for the request’s runtime state, e.g., the KV cache occupies the same GPU as the model. This limits the number of simultaneously served requests. Figure 1 shows the memory consumption for fine-tuning a single rank-8 LoRA adapter. Even at moderate sequence lengths, the runtime state requires GBs of GPU memory.
@@
Figure 1: GPU memory consumption for fine-tuning of a single rank-8 LoRA adapter for GPT2-XL, Llama2-7B, and Granite-20B.
@@
Figure 2: GPU memory layout with baseline (left) and Symbiosis (right).
@@
We address the above problems in Symbiosis. Symbiosis shares the common model parameters across the clients for inference and/or fine-tuning jobs. We refer to the shared fraction of the model as base model. Since the base model constitutes a majority of the parameters, sharing significantly reduces the GPU memory requirement and improves GPU utilization. Moreover, we transparently separate the execution of the base model (in base executor) from that of the job specific components, namely, attention and adapter layers, e.g., LoRA, (in client). This split execution results in the following benefits. (1) The separation of execution allows a single base executor to serve multiple inference or fine-tuning clients. (2) The base executor can be placed independently from its clients, e.g., the base executor can span across multiple GPUs, or the clients can be placed on different GPUs or CPUs. This is important because the clients’ resource consumption pattern differs from that of the base executor. The client contains request’s runtime state, namely, KV cache or optimizer state, which grows linearly with each new generated token (for inference) or each new pass (for fine-tuning). In contrast, the base executor’s resource consumption is proportional to the tokens processed at a given point in time. Moreover, the client is computationally less intensive, thus can be hosted on less powerful GPUs or CPUs with little performance degradation. (3) Each user can independently configure its client for an inference or fine-tuning job with a preferred PEFT method while sharing the base model instance. Users opt for different PEFT methods for their different user-case specific strengths, inference overheads and fine-tuning resource requirements [20, 22, 24, 15].
@@
  • Symbiosis proposes a general purpose framework to transparently share base model(s) across multiple inference and fine-tuning jobs located across GPUs on same or different nodes.
  • Symbiosis presents a split-execution technique, which transparently alters the model structure to decouple the execution of base model and client. This means Symbiosis works for a variety of model architectures (e.g., Llama, GPT) and a variety of PEFT methods (e.g., LoRA, IA3) without requiring any changes to model code.
  • Symbiosis presents an opportunistic batching technique to batch the inference and fine-tuning requests from different clients at the base executor, thereby improving computational efficiency. Even when the requests are of different token lengths, base executor can batch them together without requiring any padding for shorter requests.
  • Symbiosis presents a technique to preserve the privacy of model activations communicated across the clients and base executor. This enables the privacy sensitive clients to use an un-trusted base model service providers without having to expose the activations.
We demonstrate the the following use-cases in Symbiosis.
@@
Symbiosis enables the cloud providers to serve a base model instance to multiple users for inference and fine-tuning jobs. This use case achieves the following goals. First, it reduces the model serving cost for users by reducing their burden of hosting an entire model. The users can share the common base model with other users. Second, it simplifies the service provider’s resource management. It offloads the responsibility of fulfilling the job specific resource demand to users. For instance, the sequence length of a request or an optimizer selected for fine-tuning determines the GPU memory requirement. The service providers are isolated from such requirements. They only need to provision the base executor resources to fulfill the target token processing rate for the given base model, while the per-token resource requirement remains constant irrespective of the client-side configurations.
@@
Upgrading hardware with the release of new GPUs results in nodes containing different generations of GPUs. Lack of consideration for their different capabilities degrades the workload performance. Symbiosis proposes a better way leverage their different capabilities. It decouples the execution of computationally heavy base model from lighter clients. This allows the clients to be placed on less powerful devices without significant impact on inference or fine-tuning performance. Moreover, Symbiosis also enables execution of clients on CPUs. This is particularly useful for jobs with large KV cache or longer sequence lengths that require 100s of GBs of memory. Longer contexts are common in Retrieval Augmented Generation (RAG) [16], where an additional context is retrieved from a knowledge source and appended to the original prompt. Leveraging CPU memory for such jobs reduces fine-tuning and inference cost. We demonstrate in evaluation that for 64K context length, Symbiosis performs CPU-GPU inference and achieves 33% speed up compared the baseline.
@@
With Symbiosis, multiple customers can leverage a third-party base model service to fine-tune their models. Even though the popular base model parameters are publicly available, the customers do not wish to expose their adapter parameters (which may be trained on customer’s confidential data) and activations to the third-party base model service provider. The challenge of separating the adapter parameters is solved through decoupled execution, where a tenant can host the tenant specific computation in a secure environment. Moreover, Symbiosis also avoids exposing the activations that are communicated across client and base executor in order to protect the adapter parameters against model extraction attacks [21, 34, 12]. These attack observe the model activations to infer the model parameters.
2 Related Work
2.1 Simultaneous Fine Tuning and Inference
@@
Most commonly used PEFT fine-tuning systems (like PyTorch, HF-Trainer [36]) target single adapter fine-tuning where multi-adapter fine-tuning can only be achieved via multiple isolated tuning jobs. MixLoRA [17] supports fine-tuning of multiple LoRA adapters for multiple experts in Mixture of Expert model. In contrast, Symbiosis proposes a general purpose system. FlexLLM [25] enables sharing of the base model across multiple fine-tuning jobs. However, Symbiosis offers better as-a-service model because it isolates the clients’ variable GPU memory demand (from growing KV cache or optimizer state) from each other and from the base executor. We demonstrate that such isolation allows better use of heterogeneous resources and enables privacy of client activations and parameters. mLoRA [39, 43] can simultaneously fine-tune multiple LoRA adapters by sharing the base model across the trainers. However, it only supports variants of LoRA fine-tuning methods. Whereas, Symbiosis allows fine-tuning of wider variety PEFT methods, e.g., IA3, P-tuning, Prefix-tuning etc. Second, unlike mLoRA, Symbiosis does not need to modify the model or user code. Therefore, it can work with all the supported models in HuggingFace Transformers library [36] out-of-the-box. Third, mLoRA only supports pipeline parallelism to scale out a model that does not fit on a single GPU. Instead, we support Fully Sharded Data Parallel (FSDP) fine-tuning. FSDP squeezes out bubbles by overlapping computation and communication and improves training efficiency. Such bubbles are common in pipeline parallelism [18, 42]. Finally, a key difference between mLoRA and Symbiosis is that in Symbiosis, each client is independent and is a driver of its training or inference. Because of this design principle each client can independently decide its training speed, select different PEFT fine-tuning method, and determine its physical location.
@@
For inference, vLLM [14, 31], Punica [6], LoRAX [9], Caraserve [19] enable sharing of a base model across different adapters and reduce memory consumption. Moreover, their heterogeneous batching allows LoRA adapters with different characteristics, e.g., rank, to be batched together. dLoRA [37] dynamically switches between merged and un-merged adapter to reduce request latency. However, during inference, different adapters execute as a part the same process as the base model. This prevents isolation of the tenant specific data, specifically the adapter parameters, KV cache and activations. This means the data cannot be maintained privately and client’s resources cannot be managed independently. Moreover, the above platforms execute multiple adapters in lock step. Which means the clients with shorter sequences or smaller batches need to wait for the execution of the clients with longer sequences or larger batches.
2.2 Offloading Layers and Attention
Transformers library [35] natively supports offloading of KV cache to CPU to accommodate large KV cache [1]. Attention offloading is also well researched in LLMs. The attention is memory bound as opposed to the linear layers. FlexGen [32], Lamina [7], LayerKV [38] and DeepSpeed-Inference [13] offload attention computation to CPU while offloading the KV caches to host memory and storage. LMDeploy [8], CachedAttention [10] offload KV caches to host memory and storage for multi-turn conversations. InstInfer [26] offloads attention to computational storage and uses its internal bandwidth and computational capability to improve inference. For training and fine-tuning, DeepSpeed [5, 29, 30] offloads model parameters to host memory and fetches them on-demand. L2L [28] only swaps the necessary model parameters and state, layer by layer, from host memory into the GPU memory to enable constant resource training.
In contrast to the above techniques, the users of Symbiosis can independently decide whether to offload the KV cache and attention calculation to CPU (or another GPU) based on their workload.Moreover, Symbiosis presents a unique technique where users can leverage both faster GPU and memory-rich CPU for processing extremely long sequences.
2.3 Efficient Batching for Performance
@@
Efficient batching is well explored for LLMs [11, 40, 27]. Orca [40] proposes continuous batching, which instead of waiting for all requests in a batch to complete, batches new requests along with the ongoing requests. This reduces wasted computation on differently sized requests and improves throughput. Several research works also leverage the different resource consumption and performance characteristics of prefill and generation phases. Specifically, the prefill phase can saturate GPUs at smaller batches, whereas the generation phase performs better at larger batches. Sarathi [3] proposes chunked prefill to split the prefill requests into smaller chunks and the remaining slots are filled with generation phase requests. This leads to better utilization of GPUs and improves inference throughput. Splitwise [27] splits the prefill and generation phases on different machines and allows phase specific resource management. It uses cluster-level scheduling to direct a request to a machine based on its phase. It migrates KV cache from the prefill machine to the generation machine as the requests enters generation.
Symbiosis fundamentally differs from the above works in that its base model layers serve a mix of inference (prefill or generation) and fine-tuning requests. Because of their different execution speeds and performance goals, in Symbiosis, the requests batched for the execution of the first layer may not be batched together for the execution of the subsequent layers. Therefore, we employ per-layers batching policies that honor low latency and improve system throughput.
3 Design and Implementation
3.1 Design Goals
  1. 1.

### v2

@@ Abstract.
Parameter-efficient fine-tuning (PEFT) allows model builders to capture the task-specific parameters into adapters, which are a fraction of the size of the original base model. Popularity of PEFT technique for fine-tuning has led to the creation of a large number of adapters for popular Large Language Models (LLMs). However, existing frameworks fall short in supporting inference or fine-tuning with multiple adapters in the following ways. 1) For fine-tuning, each job needs to deploy its dedicated base model instance, which results in excessive GPU memory consumption and poor GPU utilization. 2) While popular inference platforms can serve multiple PEFT adapters, they do not allow independent resource management or mixing of different PEFT methods. 3) They cannot make effective use of heterogeneous accelerators. 4) They do not provide privacy to users who may not wish to expose their fine-tuned parameters to service providers. In Symbiosis, we address the above problems by enabling the as-a-service deployment of the base model. The base model layers can be shared across multiple inference or fine-tuning processes. Our split-execution technique decouples the execution of client-specific adapters and layers from the frozen base model layers offering them flexibility to manage their resources, to select their fine-tuning method, to achieve their performance goals. Our approach is transparent to models and works out-of-the-box for most models in the transformers library. Our evaluation on Llama2-13B shows that compared to baseline, Symbiosis can fine-tune 4X more adapters on the same set of GPUs in the same amount of time.
conference: ; ;
1. Introduction
@@
Parameter Efficient Fine Tuning (PEFT) (Mangrulkar et al., 2022) is a popular method to fine-tune the existing foundation models with various fine-tuning methods, e.g., LoRA, IA3, and AdaLoRA. The adapters created using PEFT are a fraction of the size of the base model used, often consuming only 10s of MBs of accelerator memory, as opposed to the base model that consumes several Gigabytes. The resulting efficiency and low cost in fine-tuning adapters have led to the creation of hundreds of adapters for many popular models, such as Llama, Gemma, StarCoder, etc. For instance, an LLM agentic system may fine-tune and serve several adapters that address different use cases. Or, they may want to evaluate a range of adapter hyperparameters, e.g., LoRA rank, alpha, etc. to reach the best possible model accuracy.
However, the existing platforms (Wolf et al., 2020a) that support simultaneous fine-tuning of multiple PEFT adapters tend to underutilize GPU’s computational capability (Patel et al., 2024; Sheng et al., 2024). Users must launch separate jobs for different adapters, each requiring a dedicated model instance. The model instances occupy GPU memory while not fully exploiting its computational capability, especially when the jobs do not receive enough requests. Recent works (Chen et al., 2023; Sheng et al., 2024; Zhengmao et al., 2023; Oliaro et al., 2025) address the problem by sharing the model instance across inference or fine-tuning jobs to reduce its memory footprint and increase the computational density. However, when multiple jobs share the same GPUs, their runtime state (e.g., KV cache, optimizer state) competes with model instance and other jobs for GPU resources (Figure 1). In such systems, sharing GPUs across different jobs entails the burden of having to deal with their dynamic GPU memory demands from either varying rate of requests or context lengths.
@@
Figure 1. GPU memory consumption for fine-tuning of a single rank-8 LoRA adapter for GPT2-XL, Llama2-7B, and Granite-20B. Runtime state requires GBs of GPU memory, especially at larger sequence lengths.
@@
Figure 2. GPU memory layout with baseline (left) and Symbiosis (right).
@@
We address the above problems in Symbiosis. Symbiosis shares the common model parameters across the clients for inference and/or fine-tuning jobs. The shared fraction of the model is referred to as base model. We transparently separate the execution of the base model (in base executor) from that of the job specific components, namely, attention and adapter layers, e.g., LoRA, (in client). This split execution offers the following benefits. (1) The separation of execution allows a single base executor to serve multiple inference or fine-tuning clients, where the base executor can be placed in one or more GPUs independently from its clients. This means the base executor can span across multiple GPUs and the clients can be either co-located with the base executor, placed on different GPUs, or even placed on a different node. This is important because the clients’ resource consumption pattern differs from that of the base executor. The client contains request’s runtime state, namely, KV cache or optimizer state, which grows linearly with each new generated token (for inference) or each new pass (for fine-tuning). In contrast, the base executor is stateless. Its resource consumption directly reflects the number of tokens being actively processed at any given moment. Therefore, by separating the base executor from the client, the client can be placed independently and remains insulated from runtime state fluctuations caused by other clients. (2) The client is computationally less intensive, thus they can be hosted on less powerful GPUs or CPUs with little performance degradation. (3) Each user can independently configure its client for an inference or fine-tuning job with a preferred PEFT method while sharing the base model instance. Users opt for different PEFT methods for their different use-case specific strengths, inference overheads, and fine-tuning resource requirements (Lialin et al., 2024; Liu et al., 2022; Nikdan et al., 2024; LeewayHertz, 2025). (4) Different clients using the common base executor need not be in lockstep for the execution of base model layers. Each client can drive its execution at a different rate while opportunistically sharing the execution of the base model with other clients at a layer granularity.
@@
  • Symbiosis proposes a general purpose framework to share base model(s) across multiple inference and fine-tuning jobs located across GPUs on same or different nodes.
  • Symbiosis presents a split-execution technique, which identifies the base model and client-specific parts of the model structure and transparently decouples them. This means Symbiosis works for a variety of model architectures (e.g., Llama, GPT) and a variety of PEFT methods (e.g., LoRA, IA3) without requiring any changes to model code. We demonstrate this with 4 different model architectures.
  • Symbiosis presents an opportunistic batching technique to batch the inference and fine-tuning requests from different clients at the base executor, thereby improving computational efficiency. Even when the requests are of different token lengths, base executor can batch them together without requiring any padding for shorter requests, avoiding the computational overhead of padding.
  • Symbiosis presents a technique to preserve the privacy of model activations communicated across the clients and base executor. This enables the privacy sensitive clients to use an un-trusted base model service provider without having to expose the activations.
We demonstrate the following use-cases in Symbiosis.
@@
Symbiosis enables the cloud providers to serve a base model instance to multiple users for inference and fine-tuning jobs. This use case achieves the following goals. First, it reduces the model serving cost for users by reducing their burden of hosting an entire model. The users can share the common base model with other users. Second, it simplifies the service provider’s resource management. It offloads the responsibility of fulfilling the job specific resource demand to users. For instance, a request’s context length or an optimizer selected for a fine-tuning job determines the client’s GPU memory requirement. The service providers are isolated from such requirements. They only need to provision the base executor resources to fulfill the expected token processing rate for a given base model, which is made easier because the per-token resource requirement remains constant irrespective of the client-side configurations.
@@
Hardware upgrade with the release of new GPUs is often staggered, which results in nodes in a cluster containing different generations of GPUs. Lack of consideration for their different capabilities degrades the workload performance. Symbiosis proposes a better way to leverage their different capabilities. It decouples the execution of computationally heavy base model from the comparatively lighter clients. This allows the base model to always use the more powerful devices with the clients being potentially placed on less powerful devices, without significant impact on inference or fine-tuning performance. Moreover, Symbiosis also enables execution of clients on CPUs. This is particularly useful for jobs with large KV cache (e.g., with longer sequence lengths) that require 100s of GBs of memory. Longer contexts are common in Retrieval Augmented Generation (RAG) (Lewis et al., 2020), where an additional context is retrieved from a knowledge source and appended to the original prompt. Leveraging the larger CPU memory for such jobs reduces fine-tuning and inference cost. We demonstrate in the evaluation section that for 64K context length, Symbiosis performs CPU-GPU inference and achieves 33% speed up compared to the GPU-only baseline.
@@
With Symbiosis, multiple customers can safely leverage a third-party base model service to fine-tune their models. Even though the popular base model parameters are publicly available, customers may not wish to expose their adapter parameters (which may be trained on customers’ confidential data) and activations to a third-party base model service provider. The challenge of separating the adapter parameters is naturally addressed by Symbiosis’s decoupled execution. This allows a tenant to host the tenant-specific computation in a secure environment. Moreover, Symbiosis also provides a mechanism that avoids exposing the activations that are communicated across client and base executor in order to protect the adapter parameters against model extraction attacks (Liang et al., 2023; Wang et al., 2025; Gong et al., 2020). These attack observe the model activations to infer the model parameters.
2. Related Work
2.1. Model Sharing in Fine Tuning and Inference
Fine-tuning
@@ 2.1. Model Sharing in Fine Tuning and Inference
Most commonly used PEFT fine-tuning systems (like PyTorch, HF-Trainer (Wolf et al., 2020b)) target single adapter fine-tuning where multi-adapter fine-tuning can only be achieved via multiple isolated tuning jobs. MixLoRA (Li et al., 2024) supports fine-tuning of multiple LoRA adapters for multiple experts in Mixture of Expert model. In contrast, Symbiosis proposes a general purpose system. FlexLLM (Oliaro et al., 2025) enables sharing of the base model across multiple fine-tuning jobs. However, Symbiosis offers better as-a-service model because it isolates the clients’ variable GPU memory demand (from growing KV cache or optimizer state) from each other and from the base executor. We demonstrate that such isolation allows better use of heterogeneous resources and enables privacy of client activations and parameters. mLoRA (Ye et al., 2023; Zhengmao et al., 2023) can simultaneously fine-tune multiple LoRA adapters by sharing the base model across the trainers. However, it only supports variants of LoRA fine-tuning methods. Whereas, Symbiosis allows fine-tuning of wider variety PEFT methods, e.g., IA3, P-tuning, Prefix-tuning etc. Second, unlike mLoRA, Symbiosis does not need to modify the model or user code. Therefore, Symbiosis can work with all the supported models in HuggingFace Transformers library (Wolf et al., 2020b) out-of-the-box. Finally, a key difference between mLoRA and Symbiosis is that in Symbiosis, each client is independent and is a driver of its training or inference. Because of this design principle each client can independently decide its training speed, select different PEFT fine-tuning method, and determine its physical location.
Inference
@@
For inference, vLLM (Kwon et al., 2023; Sheng et al., 2024), Punica (Chen et al., 2023), LoRAX (Contributors, 2024), Caraserve (Li et al., [n. d.]) enable sharing of a base model across different adapters and reduce memory consumption. Moreover, their heterogeneous batching allows LoRA adapters with different characteristics, e.g., rank, to be batched together. dLoRA (Wu et al., 2024b) dynamically switches between merged and un-merged adapter to reduce request latency. However, during inference in these systems, different adapters execute as a part the same process as the base model. This means, the client-specific data, i.e., adapter parameters, KV cache and activations, cannot be shielded from being accessed by the base model service provider. Also, clients cannot isolate themselves from the changing resource demands (e.g., increasing KV cache) of other clients. Moreover, the above platforms execute multiple adapters in lock step. Which means the clients with shorter sequences or smaller batches need to wait for the execution of the clients with longer sequences or larger batches at each layer. Symbiosis’s split execution addresses these challenges by isolating the clients from from each other and only opportunistically sharing the execution of base model layers.
2.2. Offloading KV Cache and Attention
Transformers (Wolf et al., 2020a), LMDeploy (Contributors, 2023), CachedAttention (Gao et al., 2024), Symphony (Agarwal et al., 2024), DeepSpeed (Aminabadi et al., 2022; Rajbhandari et al., 2022, 2021) offload KV caches to host memory and storage to accommodate more requests or to improve the efficiency of multi-turn conversations. Splitwise (Patel et al., 2024), Mooncake (Qin et al., 2025), DistServe (Zhong et al., 2024) split the prefill and generation phases on different GPUs or machines and allows phase specific resource management. They transfer KV cache from the prefill machine to the generation machine as the requests enters generation. Symbiosis leverages the OffloadCache (off, 2024) feature of Transformers to offload the KV cache to host memory after prefill. Symbiosis not only uses the host memory to store very large KV cache but also uses CPUs to execute clients and perform heterogeneous CPU-GPU inference for token generation. Moreover, unlike Symbiosis, the goals of these systems are not to provide client control and isolation. In these approaches, the KV cache of multiple clients is handled in the same manner.
Attention offloading is also well researched in LLMs. The attention computation is memory bound as opposed to the linear layers that are compute bound, thus it can better leverage non-GPU accelerators. FlexGen (Sheng et al., 2023), Lamina (Chen et al., 2024), LayerKV (Xiong et al., 2024) offload attention computation to CPU while offloading the KV caches to host memory and storage. InstInfer (Pan et al., 2024) offloads attention to computational storage and uses its internal bandwidth and computational capability to improve inference. However, the above approaches are not transparent to model architectures, thus are not generally applicable.
Infinigen (Lee et al., 2024), H2O (Zhang et al., 2023), DuoAttention (Xiao et al., 2024) perform partial attention by focusing on important KV entries, while offloading less important entries to CPU memory. This allows them to support longer contexts while reducing GPU memory requirement. While Symbiosis can also benefit from reduced attention computation on GPU, especially when the client is located on CPU, we do not alter the default attention logic of the model provided in Transformers library.
2.3. Efficient Batching for Performance
@@
Efficient batching is well explored for LLMs (Gao et al., 2018; Yu et al., 2022; Patel et al., 2024). Orca (Yu et al., 2022) proposes continuous batching for inference requests, which instead of waiting for all requests in a batch to complete, batches new requests along with the ongoing requests. This reduces wasted computation on differently sized requests and improves throughput. Several research works also leverage the different resource consumption and performance characteristics of prefill and generation phases. Specifically, the prefill phase can saturate GPUs at smaller batches, whereas the generation phase performs better at larger batches. Sarathi (Agrawal et al., 2023) proposes chunked prefill to split the prefill requests into smaller chunks and the remaining slots are filled with generation phase requests. This leads to better utilization of GPUs and improves inference throughput. Symbiosis fundamentally differs from the above works in that its base model layers serve a mix of inference (prefill or generation) and fine-tuning requests. Because of their different execution speeds and performance goals, in Symbiosis, the requests batched for the execution of the first layer need not be batched together for the execution of the subsequent layers. Therefore, we employ per-layers batching policies that honor low latency and improve system throughput.
3. Design and Implementation
3.1. Design Goals
  1. (1)